OpenStudy (anonymous):

In a recent year the world production of weat was about 5.2 x 10^8 metric tons. Canada produced 1/25 of this wheat. About how much wheat did Canada produce?

OpenStudy (anonymous):

5.2 × 10^8 x 1/25 = 5.2/25 x 10^8 = 0.208x10^8 = 2.08x10^7

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

can u finish the rest

OpenStudy (anonymous):

You are right Thanks!

OpenStudy (anonymous):

The rest of what?

OpenStudy (anonymous):

2.08x10^7

OpenStudy (anonymous):

Yes I submited my quiz and you were right.

OpenStudy (anonymous):

oh ok i wasnt sure if you had to completely solve it